# Vision-based Learning for Drones: A Survey

### 08 Dec 2023 v1

@@ 4 Vision-based Control
However, physical attacks from adversaries, such as birds and smart enemy drones, are based on their own respective decision-making entities, which are challenging to detect and predict by existing obstacle avoidance methods. These adversaries, such as kamikaze aggressors, are able to target, chase, and crash into the drone to cause serious accidents and massive losses. Hence, it is imperative for drones to go beyond the obstacle avoidance capability and acquire a physical attack avoidance capability, where the drone can detect physical attack adversaries and smartly outmaneuver them through advanced control strategies favoring flight agility.
4.1 Indirect Methods
Indirect methods [89, 27, 90, 91, 92, 93, 87, 13, 94, 95, 96] refer to extracting features from images or videos to generate visual odometry, depth maps, and 3D point cloud maps for drones to perform path planning based on traditional optimization algorithms (see Fig. 14). Obstacle states, such as 3D shape, position, and velocity, are detected and mapped before a maneuver is taken. Once the online maps are built or obstacles are located, the drone can generate a feasible path or take action to avoid the obstacles.
SOTA indirect methods generally divide the obstacle avoidance problem into several subtasks, namely perception, mapping and planning. On the perception side, depth maps are always required to generate corresponding actions or feasible trajectories. A depth map is a grey-level level/color image that can reflect the distance between the surfaces of objects from the viewpoint of the drone. Fig. 15 shows color images and corresponding depth maps from a drone’s viewpoint. The illuminance is proportional to the distance from the camera. A lighter color denotes a nearer surface, and darker areas mean further surfaces. A depth map provides the necessary distance information for drones to make decisions to avoid static and dynamic obstacles. Currently, off-the-shelf RGB-D cameras, such as the Intel RealSense depth camera D415, the ZED 2 stereo camera, and the Structure Core depth camera, are widely used for drone applications. Therefore, traditional obstacle avoidance methods can treat depth information as a direct input. However, for physical attack avoidance, monocular depth estimation is required to achieve omnidirectional depth perception, which is a challenge to address with existing methods.
On the mapping side, a point cloud map representing a set of data points in a 3D space is commonly generated. Each point has its own Cartesian coordinates and can be used to represent a 3D shape or an object. A 3D point cloud map is not from the view of a drone but constructs a global 3D map that provides global environmental information for a drone to fly. The point-cloud map can be generated from a LIDAR scanner or many overlapped images combined with depth information. An illustration of an original scene and a point cloud map are shown in Fig. 14, where the drone can travel around without colliding with static obstacles.
Planning is a basic requirement for a vision-based drone to avoid obstacles. Within the indirect methods, planning can be further divided into two categories: one is offline methods based on high-resolution maps and pre-known position information, such as Dijkstra’s algorithm [97], A-star [98] and sequential convex optimization [99]; the other is online methods based on real-time visual perception and decision-making. Online methods can be further categorized into online path planning [87, 13, 100] and artificial potential field (APF) methods [27, 101].
Figure 16: Artificial potential field methods in drones’ obstacle avoidance [101] (a) Generated artificial potential field; (b) Flight trajectory based on the APF.
@@ 4.1 Indirect Methods
In contrast to online path planning, the artificial potential field methods require less computation resources and can well cope with dynamic obstacle avoidance using limited sensor information. The APF algorithm is one of the algorithms in the robot path planning approach that uses attractive force to achieve the objective position and repulsive force to avoid obstacles in an unknown environment [102]. Falanga et al. [27] developed an efficient and fast control strategy based on the artificial potential field method to avoid fast approaching dynamic obstacles. The obstacles in [27] are represented as repulsive fields that decay over time, and the repulsive forces are generated from the first-order derivation of the repulsive fields at each time step. However, the repulsive forces computed only reach substantial values when the obstacle is very close, which may lead to unstable and aggressive behavior. Besides, artificial potential field methods are heuristic methods that cannot guarantee global optimization and robustness for drones. Hence, it is not ideal to adopt potential field methods to avoid adversarial physical attacks on drones.
Figure 17: Quadrotor drone flying in a wild environment with end-to-end reinforcement learning. [25] (a) Training process in the simulation platform; (b) Real flight test in the wild snow environment.
@@ 4.2 End-to-end Methods
In contrast to the indirect methods, which divide the whole obstacle avoidance task into multiple sub-tasks, such as perception, mapping, and planning, end-to-end methods [103, 104, 105, 25, 106] combine computer vision and reinforcement learning (RL) to map the visual observations to actions directly. RL [107] is a technique for mapping the state (observation) space to the action space in order to maximize a long-term return with given rewards. The learner is not explicitly told what action to carry out but must figure out which action will yield the highest reward during the exploring process. A typical RL model (see Fig. 18) features agents, the environment, reward functions, action, and state space. The policy model achieves convergent status via constant interactions between the agents and the environment, where the reward function guides the training process.
@@ 4.3 Semi-direct Methods
Method Type Perception End Control End Key Studies Description
@@
Focus on generating visual odometry, depth maps, and 3D point cloud maps for path planning.
@@
Utilize depth information for real-time visual perception and decision-making in dynamic environments.
@@
Combine deep learning for visual perception with RL for direct action response.
@@
Use deep learning for visual encoding and train the control system with expert demonstrations and online training.
@@
Extract intermediate features like relative positions or velocities and use DRL for action decisions.
@@
Utilize direct image data to obtain necessary states for obstacle avoidance using non-learning-based methods.
@@ 5.1 Single Drone Application
Single drone applications in vision-based learning primarily involve tasks like surveillance [136, 137, 138, 139], search-and-rescue operations [140, 141, 142], environmental monitoring [143, 144, 145, 146], industrial inspection [147, 148, 23] and autonomous racing [149, 150, 151]. Each field, while benefiting from the unique capabilities of drones, also presents its own set of challenges and areas for development.
@@ 5.1.1 Surveillance
In surveillance, drones are employed for continuous monitoring, playing a critical role in security and observational tasks. In [136], a real-time drone surveillance system was deployed to identify violent individuals in public areas. This real-time detection was facilitated by cloud processing of drone images, enabling the handling of the slow and memory-intensive computations required by the proposed SHDL network, while still maintaining onboard short-term navigation capabilities. In [139], a drone-based crowd surveillance system was tested based on face recognition technology. The Mobile Edge Computing (MEC) method was used to offload video data processing from the drones to achieve the goal of saving the scarce energy of the drone battery.
@@ 5.1.2 Search and Rescue
In the field of search and rescue, the work [140] presented and analyzed a generalized contour search algorithm and related this search to existing coverage searches, achieving superior search results. Since the scenarios for search and rescue missions are often extremely complex, such as collapsed houses, mountains and forests, how to improve the autonomous obstacle avoidance ability of the drones and rationally plan the traveling path of the drones are the main challenges faced at present.
@@ 5.1.3 Environmental Monitoring
For environmental monitoring, the use of drones is increasingly recognized as an efficient method for conducting large-scale site inspections. In [143], Senthilnath et al. demonstrated the application of fixed-wing and VTOL (Vertical Take-Off and Landing) drones for vegetation analysis, showcasing how these drones can be utilized for detailed crop region and tree crown mapping through spectral-spatial classification methods. Drones with cameras can also be used in grasslands for species classification [146]. The research contributed to the field by developing a comprehensive methodology for processing drone-acquired imagery and offering significant implications for environmental monitoring and management.
@@ 5.1.4 Industrial Inspection
Industrial inspection, another key application area, benefits from drones’ ability to access remote and hazardous locations. Drones were also utilized in measuring proximity among construction entities [147] to prevent a significant number of fatal and nonfatal injuries in the construction industry. The authors presented a drone-assisted visual monitoring method, which in turn can make timely intervention possible. Drones equipped with cameras can also be used in structural health inspection, especially on numerous measurement locations that are typically at very high elevations or largely inaccessible areas [148]. However, the challenge of lengthy calculation times persists, mainly due to the high computational requirements, and needs to be addressed in this work.
@@ 5.1.5 Autonomous Racing
Autonomous racing, as another significant application, mainly benefits from the advanced maneuverability and speed capabilities of single drones. Mini-drones are commonly used for implementing autonomous racing based on the characteristics of small size and high speed. In [149], a novel sensor fusion method was proposed to enable high-speed autonomous racing for mini-drones. This work also addressed issues with occasional large outliers and vision delays commonly encountered in fast drone racing. Another work [150] introduced an innovative approach to drone control, where a deep neural network (DNN) was used to fuse trajectories from multiple controllers. In the latest work [15], the vision-based drone outperformed world champions in the racing task, purely relying on onboard perception and a trained neural network. However, these studies highlight challenges in generalizing the perception network to different environments and textures without extensive data augmentation.
Overall, the key challenges in these applications include limited battery life, which restricts operational time, and the need for robust obstacle avoidance in diverse environments. Additionally, single drones often face limitations in data processing capabilities, which can impact real-time decision-making and environmental adaptability. These factors require advanced solutions to ensure drones can perform effectively and reliably in a wide range of scenarios.
@@ 5.2 Multi-Drone Application
Multi-drone applications, encompassing activities such as coordinated surveying [157, 158, 153, 159], cooperative tracking [154, 160, 161], synchronized monitoring [155, 162, 163], and disaster response [164, 165, 152, 57, 156], bring the added complexity of inter-drone communication, coordination and real-time data integration. These applications leverage the combined capabilities of multiple drones to achieve greater efficiency and coverage than single drone operations.
@@ 5.2.1 Coordinated Surveying
In coordinated surveying, a monocular visual odometry algorithm was used to enable autonomous onboard control of each vehicle with cooperative localization and mapping [157]. This work addressed the challenges of coordinating and merging different maps constructed by each drone platform and what’s more, the computational bottlenecks typically associated with 3D RGB-D cooperative SLAM. Micro-air vehicles also play an outstanding role in the field of coordinated surveying. In [159], a sensor fusion scheme was proposed to improve the accuracy of localization in MAV fleets. Ensuring that each MAV effectively contributes to the overall perception of the environment poses a challenge addressed in this work.
@@ 5.2.2 Cooperative Tracking
Cooperation between multi-drones for object tracking can also improve the tracking accuracy of the model. Studies like [160] explored the development of a cooperative surveillance and tracking capability for multi-drone formation in urban environments. This study also considered the challenge of collision avoidance between drones that may exist during multi-drone cooperation. Farmani et al. [161] discussed a decentralized multi-target tracking system for drones with limited sensing capabilities. The main problem this work tries to address is the challenge of limited communication bandwidth and intermittent connectivity among UAV team members.
@@ 5.2.3 Synchronized Monitoring
In synchronized monitoring, drones simultaneously gather data from different perspectives or locations. In the study [155], Gu et al. utilized a small target detection, localization, and tracking model that relied on the fusion of multiple data sources to realize a new cooperative network platform for synchronized monitoring. This study also aimed to address the common problems of overcoming challenges in task distribution and efficiency optimization. The work [162] aimed to address the challenges of implementing deep neural network (DNN) models on drones with limited computational capabilities and to optimize the classification latency in drone collaborative inference systems. However, common issues such as collision avoidance in complex environments and signal interference in multi-drone cooperation systems are not explicitly detailed in these papers.
@@ 5.2.4 Disaster Response
Disaster response operations benefit from multiple drones through rapid assessment of affected areas and aiding in search and rescue missions. In [164], a tracking-learning-detection framework and an advanced flocking strategy were used for exploring and searching missions. Some main challenges addressed in the paper include limited communication among drones, autonomous navigation and obstacle avoidance in complex environments, and, what’s more, real-time motion planning for cooperation systems. There are also some challenges not mentioned in this work, like the effective distribution of tasks to optimize overall mission efficiency.
The primary challenges in multi-drone applications include maintaining stable and reliable communication links, collision avoidance between drones, and effective distribution of tasks to optimize the overall mission efficiency. Additionally, issues like signal interference and managing the flight paths of multiple drones simultaneously are significant hurdles. These problems demand innovative solutions to harness the full potential of multi-drone operations in various application domains.
@@ 6.7 Safety and Security
Ensuring the safety and security of drone operations is paramount, especially in densely populated or sensitive areas. This includes not only physical safety but also cybersecurity concerns[193]. The security aspect extends beyond data protection, including the resilience of drones to adversarial attacks. Such attacks could take various forms, from signal jamming to deceptive inputs aimed at misleading vision-based systems and DRL algorithms [194]. Addressing these concerns requires a multifaceted approach. Firstly, incorporating advanced cryptographic techniques ensures data integrity and secure communication. Secondly, implementing anomaly detection systems can help identify and mitigate unusual patterns indicative of adversarial interference. Moreover, improving the robustness of learning models against adversarial attacks and investigating the explainability of designed models [195, 196] are imperative. Lastly, regular updates and patches to the drone’s software, based on the latest threat intelligence, can fortify its defenses against evolving cyber threats.

### 02 Jan 2024 v2

@@ 4.1 Indirect Methods
In contrast to online path planning, the artificial potential field methods require less computation resources and can well cope with dynamic obstacle avoidance using limited sensor information. The APF algorithm is one of the algorithms in the robot path planning approach that uses attractive force to achieve the objective position and repulsive force to avoid obstacles in an unknown environment [102]. Falanga et al. [27] developed an efficient and fast control strategy based on the artificial potential field method to avoid fast approaching dynamic obstacles. The obstacles in [27] are represented as repulsive fields that decay over time, and the repulsive forces are generated from the first-order derivation of the repulsive fields at each time step. However, the repulsive forces computed only reach substantial values when the obstacle is very close, which may lead to unstable and aggressive behavior. Besides, artificial potential field methods are heuristic methods that cannot guarantee global optimization and robustness for drones. Hence, it is not ideal to adopt potential field methods to navigate through cluttered environments.
Figure 16: Artificial potential field (APF) methods in drones’ obstacle avoidance [100] (a) Generated artificial potential field; (b) Flight trajectory based on the APF.
@@ 4.2 End-to-end Methods
In contrast to the indirect methods, which divide the whole mission into multiple sub-tasks, such as perception, mapping, and planning, end-to-end methods [48, 103, 104, 28, 105] combine computer vision and reinforcement learning (RL) to map the visual observations to actions directly. RL [106] is a technique for mapping the state (observation) space to the action space in order to maximize a long-term return with given rewards. The learner is not explicitly told what action to carry out but must figure out which action will yield the highest reward during the exploring process. A typical RL model (see Fig. 18) features agents, the environment, reward functions, action, and state space. The policy model achieves convergent status via constant interactions between the agents and the environment, where the reward function guides the training process.
Figure 17: Quadrotor drone flying in a wild environment with end-to-end reinforcement learning [28]. (a) Training process in the simulation platform; (b) Real flight test in the wild snow environment.
@@ 4.3 Semi-direct Methods
Method Type Perception End Control End Key Studies Description and Applicability
@@
Focus on generating visual odometry, depth maps, and 3D point cloud maps for path planning. Suitable for safety-critical tasks with accurate models.
@@
Utilize depth information for real-time visual perception and decision-making in dynamic environments. Suitable for rapid response tasks with certain information.
@@
Combine deep learning for visual perception with RL for direct action response. Suitable for complex tasks with uncertain information.
@@
Use deep learning for visual encoding and train the control system with expert demonstrations and online training. Suitable for sample-efficient tasks with expert demonstrations.
@@
Extract intermediate features like relative positions or velocities and use DRL for action decisions. Suitable for complex tasks with high generalization requirement.
@@
Utilize direct image data to obtain necessary states for obstacle avoidance using non-learning-based methods. Suitable for tasks where robust stereo information is not available.
@@ 5.1 Single Drone Application
The versatility of single drones is increasingly recognized in a variety of challenging environments. These unmanned vehicles, with their inherent advantages, are effectively employed in critical areas such as hazardous environment detection and search and rescue operations. Single drone applications in vision-based learning primarily involve tasks like obstacle avoidance [30, 28], surveillance [133, 134, 135, 136], search-and-rescue operations [137, 138, 139], environmental monitoring [140, 141, 142, 143], industrial inspection [144, 145, 23] and autonomous racing [146, 147, 148]. Each field, while benefiting from the unique capabilities of drones, also presents its own set of challenges and areas for development.
5.1.1 Obstacle Avoidance
The development of obstacle avoidance capabilities in drones, especially for vision-based control systems, poses significant challenges. Recent studies have primarily focused on static or simple dynamic environments, where obstacle paths are predictable [26, 28, 149]. However, complex scenarios involving unpredictable physical attacks from birds or intelligent adversaries remain largely unaddressed. For instance, [26, 150, 30, 27] have explored basic dynamic obstacle avoidance but do not account for adversarial environments. To effectively handle such threats, drones require advanced features like omnidirectional visual perception and agile maneuvering capabilities. Current research, however, is limited in addressing these needs, underscoring the necessity for further development in drone technology to enhance evasion strategies against smart, unpredictable adversaries.
@@ 5.1.2 Surveillance
While drones play a pivotal role in surveillance tasks, their deployment is not without challenges. Key obstacles include managing high data processing loads and addressing the limitations of onboard computational resources. In addressing these challenges, the study by Singh et al. [133] presented a real-time drone surveillance system used to identify violent individuals in public areas. The proposed study was facilitated by cloud processing of drone images to address the challenge of slow and memory-intensive computations while still maintaining onboard short-term navigation capabilities. Additionally, in the study [136], a drone-based crowd surveillance system was tested to achieve the goal of saving scarce energy of the drone battery. This approach involved offloading video data processing from the drones by employing the Mobile Edge Computing (MEC) method. Nevertheless, while off-board processing diminishes computational demands and energy consumption, it inevitably heightens the need for data transmission. Addressing the challenge of achieving real-time surveillance in environments with limited signal connectivity is an additional critical issue that requires resolution.
@@ 5.1.3 Search and Rescue
In the field of search and rescue operations, a number of challenges exist that hinder the development of drone technology. A primary challenge faced by drones is extracting maximum useful information from limited data sources. This is crucial for improving the efficiency and success rate of these missions. Goodrich et al. [137] address this by developing a contour search algorithm designed to optimize video data analysis, enhancing the capability to identify key elements swiftly. However, incorporating temporal information into this algorithm introduces additional computational demands. These increased requirements present new challenges, such as the need for more powerful processing capabilities and potentially greater energy consumption.
@@ 5.1.4 Environmental Monitoring
A major challenge in the application of drones for environmental monitoring lies in efficiently collecting high-resolution data while navigating the constraints of battery life, flight duration, and diverse weather conditions. Addressing this, Senthilnath et al. [140] showcased the use of fixed-wing and VTOL (Vertical Take-Off and Landing) drones in vegetation analysis, focusing on the challenge of detailed mapping through spectral-spatial classification methods. In another study, Lu et al. [143] demonstrated the utility of drones for species classification in grasslands, contributing to the development of methodologies for drone-acquired imagery processing, which is crucial for environmental assessment and management. While these studies represent significant steps in drone applications for environmental monitoring, several challenges persist. Future research may need to address the problems of improving the drones’ resilience to diverse environmental conditions, and extend their operational range and duration to comprehensively cover extensive and varied landscapes.
@@ 5.1.5 Industrial Inspection
In industrial inspection, drones face key challenges like safely navigating complex environments and conducting precise measurements in the presence of various disturbances. Kim et al. [144] addressed the challenge of autonomous navigation by using drones for proximity measurement among construction entities, enhancing safety in the construction industry. Additionally, Khuc et al. [145] focused on precise structural health inspection with drones, especially in high or inaccessible locations. Despite these advancements in autonomous navigation and measurement accuracy, maintaining data accuracy and reliability in industrial settings with interference from machinery, electromagnetic fields, and physical obstacles continues to be a significant challenge, necessitating further research and development in this domain.
@@ 5.1.6 Autonomous Racing
In autonomous drone racing, the central challenge is reducing delays exist in visual information processing and decision making and enhancing the adaptability of perception networks. In [146], a novel sensor fusion method was proposed to enable high-speed autonomous racing for mini-drones. This work also addressed issues with occasional large outliers and vision delays commonly encountered in fast drone racing. Another work [147] introduced an innovative approach to drone control, where a deep neural network (DNN) was used to fuse trajectories from multiple controllers. In the latest work [15], the vision-based drone outperformed world champions in the racing task, purely relying on onboard perception and a trained neural network. The primary challenge in autonomous drone racing, as identified in these studies, lies in the need for improved adaptability of perception networks to various environments and textures, which is crucial for the high-speed demands of the sport.
Overall, the primary challenges in single drone applications include limited battery life, which restricts operational duration, and the need for effective obstacle avoidance in varied environments. Additionally, limitations in data processing capabilities affect real-time decision-making and adaptability. Advanced technological solutions are essential to overcome these challenges, ensuring that single drones can operate efficiently and reliably in diverse scenarios and paving the way for future innovations.
@@ 5.2 Multi-Drone Application
While single drones offer convenience, their limited monitoring range has prompted interest in multi-drone collaboration. This approach seeks to overcome range limitations by leveraging the collective capabilities of multiple drones for broader, more efficient operations. Multi-drone applications, encompassing activities such as coordinated surveying [156, 157, 152, 158], cooperative tracking [153, 159, 160], synchronized monitoring [154, 161, 162], and disaster response [163, 164, 151, 56, 155], bring the added complexity of inter-drone communication, coordination and real-time data integration. These applications leverage the combined capabilities of multiple drones to achieve greater efficiency and coverage than single drone operations.
@@ 5.2.1 Coordinated Surveying
In coordinated surveying, several challenges are prominent: merging diverse data from individual drones, and addressing computational demands in cooperative process. These challenges were tackled by some works. In [156], a monocular visual odometry algorithm was used to enable autonomous onboard control with cooperative localization and mapping. This work addressed the challenges of coordinating and merging different maps constructed by each drone platform and what’s more, the computational bottlenecks typically associated with 3D RGB-D cooperative SLAM. Micro-air vehicles also play an outstanding role in the field of coordinated surveying. Similarly, in [158], a sensor fusion scheme was proposed to improve the accuracy of localization in MAV fleets. Ensuring that each MAV effectively contributes to the overall perception of the environment poses a challenge addressed in this work. Both methods depend on effective collaboration and communication among multiple agents. However, maintaining stable communication between drones remains a critical and unresolved issue in multi-drone operations.
@@ 5.2.2 Cooperative Tracking
In the field of multi-drone object tracking, navigating complex operational environments and overcoming limited communication bandwidth are significant challenges. These topics have also garnered significant research interest. In the study [159], researchers developed a system for cooperative surveillance and tracking in urban settings, specifically tackling the issue of collision avoidance among drones. Additionally, another work by Farmani et al. [160] explores a decentralized tracking system for drones, focusing on overcoming limited communication bandwidth and intermittent connectivity challenges. However, a persistent difficulty in this field is navigating complex external environments. Effective path planning and avoiding obstacles during multi-drone operations remain crucial challenges that require ongoing attention and innovation.
@@ 5.2.3 Synchronized Monitoring
In synchronized monitoring missions with multiple drones, the focus is on how to effectively allocate tasks among drones and improve overall mission efficiency, as well as overcoming computational limitations. Gu et al. in [154] developed a small target detection and tracking model using data fusion, establishing a cooperative network for synchronized monitoring. This study also addresses task distribution challenges and efficiency optimization. Moreover, [161] explores implementing deep neural network (DNN) models on computationally limited drones, focusing on reducing classification latency in collaborative drone systems. However, issues like collision avoidance in complex environments and signal interference in multi-drone systems are not comprehensively addressed in these studies.
@@ 5.2.4 Disaster Response
In the domain of multi-drone systems for disaster response, challenges such as autonomous navigation, communication limitations, and real-time decision-making are paramount. To address these problems, Tang et al. [163] introduced a tracking-learning-detection framework with an advanced flocking strategy for exploration and search missions. However, the study does not examine the crucial aspect of task distribution and optimization for enhancing mission efficiency in multi-drone disaster response. This shortcoming points to an essential area for further research, as effective task management is key to utilizing the full capabilities of multi-drone systems, particularly in the dynamic and urgent context of disaster scenarios.
The primary challenges in multi-drone applications include maintaining stable and reliable communication links, collision avoidance between drones, and effective distribution of tasks to optimize the overall mission efficiency. Additionally, issues like signal interference and managing the flight paths of multiple drones simultaneously are significant hurdles. Addressing these challenges is vital for unlocking the full potential of multi-drone systems, paving the way for advancements in various application domains.
@@ 6.7 Safety and Security
Ensuring the safety and security of drone operations is paramount, especially in densely populated or sensitive areas. This includes not only physical safety but also cybersecurity concerns[192, 193]. The security aspect extends beyond data protection, including the resilience of drones to adversarial attacks. Such attacks could take various forms, from signal jamming to deceptive inputs aimed at misleading vision-based systems and DRL algorithms [194]. Addressing these concerns requires a multifaceted approach. Firstly, incorporating advanced cryptographic techniques ensures data integrity and secure communication. Secondly, implementing anomaly detection systems can help identify and mitigate unusual patterns indicative of adversarial interference. Moreover, improving the robustness of learning models against adversarial attacks and investigating the explainability of designed models [195, 196] are imperative. Lastly, regular updates and patches to the drone’s software, based on the latest threat intelligence, can fortify its defenses against evolving cyber threats.